# Sequential model confidence sets

### v3

@@ 1 Introduction
The model confidence set (MCS) proposed by Hansen et al. (2011) provides a promising solution to this problem. It departs from standard practice where just a single model is selected according to some appropriate loss and nothing is said about the uncertainty associated with this selection. The MCS takes this uncertainty into account by reducing the original set of available models to a smaller set of flexible size that contains the best models with a given level of confidence. Model confidence sets are of great importance in applications without an obvious benchmark, may be easily constructed by the MCS algorithm, and are widely applied in the econometrics community; see, e.g., Weron (2014), Corbet et al. (2019) or Masini et al. (2023) amongst others.
@@ 2 Sequential model confidence sets
Our leading example of the above setting is the following. Let $1,\dots,m$ index different forecasters which all sequentially issue predictive distributions or point forecasts $(f_{1,t})_{t\in\mathbb{N}},\dots,(f_{m,t})_{t\in\mathbb{N}}$ for some unknown quantity $(Y_{t})_{t\in\mathbb{N}}$, adapted to $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. For each $i=1,\dots,m$ and $t\in\mathbb{N}$, the forecast $f_{i,t}$, which refers to the outcome $Y_{t}$, is predictable, meaning that it is only based on information until time $t-1$. To assess the forecasters, one should employ proper scoring rules or consistent scoring functions, see Supplement A for further details. Then, the time series $(L_{i,t})_{t\in\mathbb{N}}$ emerges by applying a proper scoring rule (consistent scoring function) $L$ to the forecast-observation pairs $(f_{i,t},Y_{t})_{t\in\mathbb{N}}$, that is $L_{i,t}=L(f_{i,t},Y_{t})$, $i=1,\dots,m$, $t\in\mathbb{N}$.
@@
Let $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ be the targeted sequence of superior objects. In the case of (1) or (2), we just consider a constant sequence. For a given confidence level $\alpha\in(0,1)$, we call (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ a sequence of model confidence sets or sequential model confidence sets, for short SMCS(s), for $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}$ at level $\alpha$ if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, $\mathbb{Q}(\exists t\geq 1:\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t})\leq\alpha$, or equivalently
@@
In the paper by Hansen et al. (2011), the authors make the assumption that the (unconditional) expectation $\mathbb{E}(d_{ij,t})=\mu_{ij}$ does not depend on $t$, and they define the set of superior objects as =$\mathcal{M}^{\star}=\left\{i\in\mathcal{M}_{0}\mid\mu_{ij}\leq 0\text{ for all}j\in\mathcal{M}_{0}\right\}.$ This is in line with Diebold and Mariano (1995), who also define model (forecast) superiority with respect to the unconditional expected loss differences. For a fixed sample size, they target an asymptotic guarantee of the form (^$\mathbb{Q}(\mathcal{M}^{\star}\subseteq\widehat{\mathcal{M}})\geq 1-\alpha$ for their model confidence set (MCS) $\widehat{\mathcal{M}}$. Our method provides a stronger guarantee but it is also the first method with finite sample validity for a fixed sample size.
Model confidence sets are also related to the literature on inference on the argmin of a vector of means, see, e.g. Futschik and Pflug (1995); Zhang et al. (2024). In this strand of literature, the desired coverage guarantee is typically weaker in that $\mathbb{Q}(i\in\widehat{\mathcal{M}})\geq 1-\alpha$ should hold for each superior model $i$ but not uniformly over all models.
@@ 2.1 Discussion of the different sets of superior models
If it is reasonable to assume that some models outperform all other models at all time points (in terms of the conditional expected score differences), then we should try to find estimators for $\mathcal{M}^{\textrm{s},\star}$. We refer to this assumption as the strong hypothesis. The strong hypothesis particularly applies if we assume independent and identically distributed scores. Then, one typically studies the unconditional expected score differences $\mathbb{E}(d_{ij,t})$ independent of $t\in\mathbb{N}$, as it is done by Diebold and Mariano (1995) or Hansen et al. (2011), who assume stationary performances of the models. Another important instance where the strong hypothesis applies is if the models have nested information sets. In Section 5.1, we construct SMCSs for different Covid-19 related deaths where the assumption of nested information sets seems plausible.
@@
We sample i.i.d. standard normally distributed observations $(Y_{t})_{t=1}^{n}$ and compare $m=3$ different forecasters $i=1,2,3$ which issue median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, for $\varepsilon_{1,t}=\beta,\varepsilon_{2,t}=\gamma^{t},\varepsilon_{3,t}=\delta t,$ for $\beta,\delta>0$ and $10<\gamma<$ 1. We call forecaster $1$ constantly biased, forecaster $2$ improving and forecaster $3$ worsening. We assess them by $L(m,y)=0.5(\Phi(m)-\Phi(y))$, for $m,y\in\mathbb{R}$ and $\Phi$ the cdf of the standard normal distribution, which is a consistent scoring function for the median, see Supplement A. If we choose the parameters as given in Section 4, we have $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $550153<t<$ 550, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$. That is, whereas the worsening forecaster is initially superior, the constantly biased forecaster catches up after some time until the improving forecaster becomes the best in the end.
@@ 3.1 Sequential testing methods
A (super-)martingale with respect to a probability measure $\mathbb{Q}\in\mathfrak{B}(\Omega)$ is a sequence of $\mathbb{Q}$-integrable random variables $(X_{t})_{t\in\mathbb{N}}$, which is adapted to $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ and satisfies the condition $\mathbb{E}_{\mathbb{Q}}[X_{t}|\mathcal{F}_{t-1}]=X_{t-1}$ ($\mathbb{E}_{\mathbb{Q}}[X_{t}|\mathcal{F}_{t-1}]\leq X_{t-1}$), for all $t\in\mathbb{N}$. We call a process $(M_{t})_{t\in\mathbb{N}}$ a test (super)martingale for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ if it is a nonnegative (super)martingale with respect to any $\mathbb{Q}\in\mathcal{H}$, and has a starting value $M_{0}\leq 1$.
Test martingales allow to phrase statistical testing as a betting game. Indeed, the value of the test martingale can be interpreted as the accumulated wealth of a gambler after having bet a number of times against the null hypothesis (Shafer, 2021). For example, in a coin-toss game with probability of heads equal to $q\in(0,1)$, a player that bets 1\$ on heads in each round and receives 2\$ in case of success or 0\$ otherwise, cannot expect to gain money over time if the null hypothesis $\mathcal{H}_{0}:q\leq 1/2$ is true. Thus, large values of the test martingale give evidence against the null hypothesis.
@@
This lays the foundation for safe anytime-valid inference. By (5), test martingales yield valid sequential tests $(\psi_{t})_{t\in\mathbb{N}}$ by defining $\psi_{t}=\mathbbm{1}\{M_{t}\geq 1/\alpha\}$. Furthermore, Ville’s inequality enables the construction of time-uniform confidence sequences (Howard et al., 2021; Ramdas et al., 2023). For $\alpha\in(0,1)$, a $(1-\alpha)$-confidence sequence for a parameter sequence $(\theta_{t})_{t\in\mathbb{N}}$ in some space $\Theta$ is a sequence of sets $(C_{t})_{t\in\mathbb{N}}$ such that $\mathbb{Q}\left(\forall t\in\mathbb{N}:\theta_{t}\in C_{t}\right)\geq 1-\alpha$, for all $\mathbb{Q}\in\mathfrak{B}(\Omega).$ The definition of confidence sequences resembles that of confidence sets, with the important difference that confidence sequences provide time-uniform coverage guarantees. We will use confidence sequences to construct SMCSs for the sequence of weakly superior objects.
It has been shown that for some hypotheses $\mathcal{H}$, it is not possible to construct non-trivial test supermartingales (Ramdas et al., 2022; Henzi and Law, 2024). However, it may still be possible to find non-trivial e-processes. An e-process for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a nonnegative adapted stochastic process $(E_{t})_{t\in\mathbb{N}}$ with $\mathbb{E}_{\mathbb{Q}}(E_{\tau})\leq 1$ for all $\mathbb{Q}\in\mathcal{H}$ and all (possibly infinite) stopping times $\tau$. Equivalently, an adapted nonnegative process $(E_{t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}$ if it is upper bounded by a test supermartingale for each $\mathbb{Q}\in\mathcal{H}$. Hence, any test supermartingale is also an e-process. Importantly, Ville’s inequality continues to hold for e-processes (Ramdas et al., 2020).
@@ 3.2 SMCSs for the strongly superior and uniformly weakly superior objects
We define the strong hypotheses $\mathcal{H}^{\textrm{s}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,r}\leq 0,\forall t\in\mathbb{N}\}$, and the weak hypotheses $\mathcal{H}^{\textrm{uw}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\Delta_{ij,r}\leq 0,\forall t\in\mathbb{N}\}$ for $i,j\in\mathcal{M}_{0}$. We would like to test all these pairwise hypothesis simultaneously at each time point $t\in\mathbb{N}$. For $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, assume that, for any $i,j\in\mathcal{M}_{0}$, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for the hypothesis $\mathcal{H}^{\bullet}_{ij}$. Then, for any $i\in\mathcal{M}_{0}$, the arithmetic mean $E_{i\cdot,t}=1/(m-1)\sum_{j\neq i}E_{ij,t}$ is an e-process for the intersection hypothesis $\mathcal{H}^{\bullet}_{i\cdot}=\cap_{j\neq i}\mathcal{H}^{\bullet}_{ij}$. We adjust the e-processes $(E_{1\cdot,t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}}$ for multiple testing using the closure principle with the arithmetic mean as $e$-merging function, that is,
@@
see Supplement C for further details. As shown in Vovk and Wang (2021, Proposition 3.1.), the arithmetic mean essentially dominates all other symmetric e-merging functions. Algorithm 1 of Vovk and Wang (2021) allows for an efficient computation of the minimum at (6). Finally, for some significance level $\alpha\in(0,1)$, we define
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}^{\star}<1/\alpha\right\},\quad t\in\mathbb{N}.$ (7)
@@
For any $\alpha\in(0,1)$, the sequence (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (7) is an SMCS at level $\alpha$ for $\mathcal{M}^{\bullet,\star}$, $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, and so is its running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$.
Theorem 3.1 works since we control the family-wise error rate of the tests $\psi_{i\cdot,t}^{\star}$ over all models $i$. Family-wise error rate control relies on the closure principle (Marcus et al., 1976) and some e-merging function, see Supplement C. Alternatively, one could adjust the pairwise e-processes $E_{ij,t}$ directly to obtain adjusted e-processes for the pairwise hypotheses $\mathcal{H}^{\bullet}_{ij}$ and exclude the model $i$ from the SMCS at $t\in\mathbb{N}$ if $E^{\star}_{ij,t}\geq 1/\alpha$ for some $j\neq i$. However, the corresponding tests can be shown to be uniformly less powerful, see Supplement C.
@@
Except from the family-wise error rate, the false discovery rate (FDR) is arguably the most commonly used criterion in multiple testing (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001). Recently, Wang and Ramdas (2022) studied FDR control for e-values. In principle, SMCSs could be constructed based on FDR control. However, this fundamentally changes the type of validity we are obtaining for ^$\widehat{\mathcal{M}}_{t}$. It allows us to bound the fraction of wrongly rejected models by the total number of rejected models at all time points, which is different from our target coverage guarantee at (4). We give some further details in Supplement D.

### v4

@@ 1 Introduction
The model confidence set (MCS) proposed by Hansen et al. (2011) provides a promising solution to this problem. It departs from standard practice where just a single model is selected according to some appropriate loss and nothing is said about the uncertainty associated with this selection. The MCS takes this uncertainty into account by reducing the original set of available models to a smaller set of flexible size that contains the best models with a given level of confidence. Model confidence sets are of great importance in applications without an obvious benchmark, may be easily constructed by the MCS algorithm, and are widely applied in the econometrics community; see, e.g., Nan (2009) and Peng et al. (2018) amongst others.
@@ 2 Sequential model confidence sets
Our leading example of the above setting is the following. Let $1,\dots,m$ index different forecasters which all sequentially issue predictive distributions or point forecasts $(f_{1,t})_{t\in\mathbb{N}},\dots,(f_{m,t})_{t\in\mathbb{N}}$ for some unknown quantity $(Y_{t})_{t\in\mathbb{N}}$, adapted to $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. For each $i=1,\dots,m$ and $t\in\mathbb{N}$, the forecast $f_{i,t}$, which refers to the outcome $Y_{t}$, is predictable, meaning that it is only based on information until time t1t-1. To assess the forecasters, one should employ proper scoring rules or consistent scoring functions, see Supplement A for further details. Then, the time series $(L_{i,t})_{t\in\mathbb{N}}$ emerges by applying a proper scoring rule (consistent scoring function) $L$ to the forecast-observation pairs $(f_{i,t},Y_{t})_{t\in\mathbb{N}}$, that is $L_{i,t}=L(f_{i,t},Y_{t})$, $i=1,\dots,m$, $t\in\mathbb{N}$.
@@
Let $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ be the targeted sequence of superior objects. In the case of (1) or (2), we just consider a constant sequence. For a given confidence level $\alpha\in(0,1)$, we call $(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ a sequence of model confidence sets or sequential model confidence sets, for short SMCS(s), for $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}$ at level $\alpha$ if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, $\mathbb{Q}(\exists t\geq 1:\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t})\leq\alpha$, or equivalently
@@
In the paper by Hansen et al. (2011), the authors make the assumption that the (unconditional) expectation $\mathbb{E}(d_{ij,t})=\mu_{ij}$ does not depend on $t$, and they define the set of superior objects as $\mathcal{M}^{\star}=\left\{i\in\mathcal{M}_{0}\mid\mu_{ij}\leq 0\text{ for all }j\in\mathcal{M}_{0}\right\}.$ This is in line with Diebold and Mariano (1995), who also define model (forecast) superiority with respect to the unconditional expected loss differences. For a fixed sample size, they target an asymptotic guarantee of the form $\mathbb{Q}(\mathcal{M}^{\star}\subseteq\widehat{\mathcal{M}})\geq 1-\alpha$ for their model confidence set (MCS) $\widehat{\mathcal{M}}$. Our method provides a stronger guarantee but it is also the first method with finite sample validity for a fixed sample size.
Model confidence sets are also related to the literature on inference on the argmin of a vector of means, see, e.g. Futschik and Pflug (1995); Zhang et al. (2024). In this strand of literature, the desired coverage guarantee is typically weaker in that $\mathbb{Q}(i\in\widehat{\mathcal{M}})\geq 1-\alpha$ should hold for each superior model $i$ but not uniformly over all models. We discuss an adaptation of our methods to this so-called marginal coverage guarantee in Supplement D.
@@ 2.1 Discussion of the different sets of superior models
If it is reasonable to assume that some models outperform all other models at all time points (in terms of the conditional expected score differences), then we should try to find estimators for $\mathcal{M}^{\textrm{s},\star}$. We refer to this assumption as the strong hypothesis. The strong hypothesis particularly applies if we assume independent and identically distributed scores. Then, one typically studies the unconditional expected score differences $\mathbb{E}(d_{ij,t})$ independent of $t\in\mathbb{N}$, as it is done by Diebold and Mariano (1995) or Hansen et al. (2011), who assume stationary performances of the models. Another important instance where the strong hypothesis is reasonable is if the models have nested information sets and use them in a nearly optimal way, see Holzmann and Eulert (2014). In Section 5.1, we construct SMCSs for different Covid-19 related deaths where the assumption of nested information sets seems plausible.
@@
We sample i.i.d. standard normally distributed observations $(Y_{t})_{t=1}^{n}$ and compare $m=3$ different forecasters $i=1,2,3$ which issue median predictions $m_{i,t}=Y_{t}+\varepsilon_{i,t}$, for $\varepsilon_{1,t}=\beta,\varepsilon_{2,t}=\gamma^{t},\varepsilon_{3,t}=\delta t,$ for $\beta,\delta>0$ and $0<\gamma<1$. We call forecaster $1$ constantly biased, forecaster $2$ improving and forecaster $3$ worsening. We assess them by $L(m,y)=0.5(\Phi(m)-\Phi(y))$, for $m,y\in\mathbb{R}$ and $\Phi$ the cdf of the standard normal distribution, which is a consistent scoring function for the median, see Supplement A. If we choose the parameters as given in Section 4, we have $\mathcal{M}^{\textrm{w},\star}_{t}=\{3\}$ for $t\leq 153$, $\mathcal{M}^{\textrm{w},\star}_{t}=\{1\}$ for $153<t<550$, and $\mathcal{M}^{\textrm{w},\star}_{t}=\{2\}$ for $t\geq 550$. That is, whereas the worsening forecaster is initially superior, the constantly biased forecaster catches up after some time until the improving forecaster becomes the best in the end.
@@ 3.1 Sequential testing methods
A (super-)martingale with respect to a probability measure $\mathbb{Q}\in\mathfrak{B}(\Omega)$ is a sequence of $\mathbb{Q}$-integrable random variables $(X_{t})_{t\in\mathbb{N}}$, which is adapted to $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ and satisfies the condition $\mathbb{E}_{\mathbb{Q}}[X_{t}|\mathcal{F}_{t-1}]=X_{t-1}$ ($\mathbb{E}_{\mathbb{Q}}[X_{t}|\mathcal{F}_{t-1}]\leq X_{t-1}$), for all $t\in\mathbb{N}$. We call a process $(M_{t})_{t\in\mathbb{N}}$ a test (super-)martingale for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ if it is a nonnegative (super)martingale with respect to any $\mathbb{Q}\in\mathcal{H}$, and has a starting value $M_{0}\leq 1$.
Test supermartingales allow to phrase statistical testing as a betting game. Indeed, the value of the test supermartingale can be interpreted as the accumulated wealth of a gambler after having bet a number of times against the null hypothesis (Shafer, 2021). For example, in a coin-toss game with probability of heads equal to $q\in(0,1)$, a player that bets 1\$ on heads in each round and receives 2\$ in case of success or 0\$ otherwise, cannot expect to gain money over time if the null hypothesis $\mathcal{H}_{0}:q\leq 1/2$ is true. Thus, large values of the test supermartingale give evidence against the null hypothesis.
@@
This lays the foundation for safe anytime-valid inference. A proof can be found in (Howard et al., 2020, Section 6.1). By (5), test supermartingales yield valid sequential tests $(\psi_{t})_{t\in\mathbb{N}}$ by defining $\psi_{t}=\mathbbm{1}\{M_{t}\geq 1/\alpha\}$. Furthermore, Ville’s inequality enables the construction of time-uniform confidence sequences (Howard et al., 2021; Ramdas et al., 2023). For $\alpha\in(0,1)$, a $(1-\alpha)$-confidence sequence for a parameter sequence $(\theta_{t})_{t\in\mathbb{N}}$ in some space $\Theta$ is a sequence of sets $(C_{t})_{t\in\mathbb{N}}$ such that $\mathbb{Q}\left(\forall t\in\mathbb{N}:\theta_{t}\in C_{t}\right)\geq 1-\alpha$, for all $\mathbb{Q}\in\mathfrak{B}(\Omega).$ The definition of confidence sequences resembles that of confidence sets, with the important difference that confidence sequences provide time-uniform coverage guarantees. We will use confidence sequences to construct SMCSs for the sequence of weakly superior objects.
It has been shown that for some hypotheses $\mathcal{H}$, it is not possible to construct non-trivial test supermartingales (Ramdas et al., 2022; Henzi and Law, 2024). However, it may still be possible to find non-trivial e-processes. An e-process for some hypothesis $\mathcal{H}\subseteq\mathfrak{B}(\Omega)$ is a nonnegative adapted stochastic process $(E_{t})_{t\in\mathbb{N}}$ with $\mathbb{E}_{\mathbb{Q}}(E_{\tau})\leq 1$ for all $\mathbb{Q}\in\mathcal{H}$ and all (possibly infinite) stopping times $\tau$. Equivalently, an adapted nonnegative process $(E_{t})_{t\in\mathbb{N}}$ is an e-process for $\mathcal{H}$ if, for each $\mathbb{Q}\in\mathcal{H}$, it is upper bounded by a test supermartingale. In other words, the upper bounding test supermartingale may be different for each $\mathbb{Q}\in\mathcal{H}$. Hence, any test supermartingale is also an e-process. Importantly, Ville’s inequality continues to hold for e-processes (Ramdas et al., 2020).
@@ 3.2 SMCSs for the strongly superior and uniformly weakly superior objects
We define the strong hypotheses $\mathcal{H}^{\textrm{s}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\mu_{ij,r}\leq 0,\forall t\in\mathbb{N}\}$, and the weak hypotheses $\mathcal{H}^{\textrm{uw}}_{ij}=\{\mathbb{Q}\in\mathfrak{B}(\Omega)\mid\Delta_{ij,r}\leq 0,\forall t\in\mathbb{N}\}$ for $i,j\in\mathcal{M}_{0}$. We would like to test all these pairwise hypothesis simultaneously at each time point $t\in\mathbb{N}$. For $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, assume that, for any $i,j\in\mathcal{M}_{0}$, $(E_{ij,t})_{t\in\mathbb{N}}$ is an e-process for the hypothesis $\mathcal{H}^{\bullet}_{ij}$. Then, for any $i\in\mathcal{M}_{0}$, the arithmetic mean $E_{i\cdot,t}=1/(m-1)\sum_{j\neq i}E_{ij,t}$ is an e-process for the intersection hypothesis $\mathcal{H}^{\bullet}_{i\cdot}=\cap_{j\neq i}\mathcal{H}^{\bullet}_{ij}$. We adjust the e-processes $(E_{1\cdot,t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}}$ for multiple testing using the closure principle (Marcus et al., 1976) with the arithmetic mean as $e$-merging function, that is,
@@
see Supplement C for further details. Importantly, each adjusted process $(E_{i\cdot,t}^{\star})_{t\in\mathbb{N}}$ is an e-process for the hypothesis $\mathcal{H}_{i\cdot}^{\bullet}$. As shown in Vovk and Wang (2021, Proposition 3.1.), the arithmetic mean essentially dominates all other symmetric e-merging functions. Algorithm 1 of Vovk and Wang (2021) allows for a computation of the minimum at (6) with computation time $O(m^{2})$. In fact, the computation time may be improved to $O(m\log m)$, see Supplement H for details. Finally, for some significance level $\alpha\in(0,1)$, we define
$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}^{\star}<1/\alpha\right\},\quad t\in\mathbb{N}.$ (7)
@@
For any $\alpha\in(0,1)$, the sequence $(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}$ defined at (7) is an SMCS at level $\alpha$ for $\mathcal{M}^{\bullet,\star}$, $\bullet\in\{\mathrm{s},\mathrm{uw}\}$, and so is its running intersection $\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$.
Theorem 3.1 works since we control the family-wise error rate of the tests $\psi_{i\cdot,t}^{\star}$ over all models $i$. Family-wise error rate control relies on the closure principle and some e-merging function, see Supplement C. Alternatively, one could adjust the pairwise e-processes $E_{ij,t}$ directly to obtain adjusted e-processes for the pairwise hypotheses $\mathcal{H}^{\bullet}_{ij}$ and exclude the model $i$ from the SMCS at $t\in\mathbb{N}$ if $E^{\star}_{ij,t}\geq 1/\alpha$ for some $j\neq i$. However, the corresponding tests can be shown to be uniformly less powerful, see Supplement C.
@@
Except from the family-wise error rate, the false discovery rate (FDR) is arguably the most commonly used criterion in multiple testing (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001). Recently, Wang and Ramdas (2022) studied FDR control for e-values. In principle, SMCSs could be constructed based on FDR control. However, this fundamentally changes the type of validity we are obtaining for $\widehat{\mathcal{M}}_{t}$. It allows us to bound the fraction of wrongly rejected models by the total number of rejected models at all time points, which is different from our target coverage guarantee at (4). We give some further details in Supplement E.
Remark 2.
Our anytime-valid methods readily extend if one is willing to weaken the simultaneous guarantee given at (4) by marginal coverage, as defined right before Section 2.1, thereby allowing to omit the adjustment step for multiple testing. The same potential gain in power may be achieved under the additional assumption that there is only one single superior object. See Supplement D for further details.